# The causal effects of modified treatment policies under network interference

### v2

@@ 3.1 Induced Modified Treatment Policies
As discussed in Section 2, identifiability can be restored despite the presence of interference by performing inference on $A^{s}$ instead of $A$ and $L^{s}$ instead of $L$. This approach is, however, incompatible with the application of MTPs: one must consider not the causal effect of $A$ under $d(\cdot;\delta)$; but rather, the effect of $A^{s}$ after intervening on the upstream exposure via $d(\cdot;\delta)$. To identify the causal effects of MTPs under interference, we introduce a novel intervention scheme—the induced MTP.
Consider applying an MTP to the SCM (4), replacing $\mathbf{A}$ with $\mathbf{A}^{d}=d(\mathbf{A},\mathbf{L};\delta)=[d(A_{i},L_{i};\delta)]_{i=1}^{n}$. Under interference, we are interested in the causal effect of $A^{s}$ on $Y$. Hence, the scientific question of interest is actually, “what if $A_{i}^{s}$ were replaced by some $A_{i}^{s\circ d}=s_{F_{i}}(d(\mathbf{A},\mathbf{L};\delta),\mathbf{L})$?” This process is illustrated in Figure 1. We call the function composition $s\circ d$ the induced MTP.
@@
Under an induced MTP, interference no longer hampers identifiability because $s_{F_{i}}(\mathbf{A},\mathbf{L})$ captures the contribution of all relevant units (that is, a given unit $i$ and its neighbors) to each $Y_{i}$. This data-adaptive parameter will converge to the population counterfactual mean as $n\rightarrow\infty$ (Hubbard2016). Use of such a parameter definition is necessary because we must condition on the single observation of the interference network at play. Hence, do note that in all theory that follows, our estimates will implicitly condition on the observed network $\mathbf{F}$. This estimand must be compatible with the network: it is interpreted as the average change in $Y$ caused by imposing the unit-level MTP $d$ on each unit in the population governed by the network $\mathbf{F}$. With $\Psi_{n}(\mathsf{P})$ identified, the population intervention effect (a contrast) may be defined by subtracting $\mathbb{E}Y$ (munozPopulationInterventionCausal2012).
@@ 3.2 Identification
If $(s_{F_{i}}(\mathbf{a},\mathbf{l}),s_{F_{i}}(\mathbf{l}))\in\text{supp}\{A_{i}^{s},L_{i}^{s}\}$, then $(s_{F_{i}}(\mathbf{a}^{d},\mathbf{l}),s_{F_{i}}(\mathbf{l}))\in\text{supp}\{A_{i}^{s},L_{i}^{s}\}$.

### v3

@@ 3.1 Induced Modified Treatment Policies
As discussed in Section 2, restoring identifiability by performing inference on $A^{s}$ and $L^{s}$ instead of $A$ and $L$ is a common strategy for handling interference. This approach is, however, not compatible with the application of MTPs: one must consider not the causal effect of $A$ under $d(\cdot;\delta)$; but rather, the effect of $A^{s}$ after intervening on the upstream exposure via $d(\cdot;\delta)$. To identify the causal effects of MTPs under interference, we introduce a novel intervention scheme—the induced MTP.
Consider applying an MTP to the SCM (4), replacing $\mathbf{A}$ with the vector $\mathbf{A}^{d}=d(\mathbf{A},\mathbf{L};\delta)=[d(A_{i},L_{i};\delta)]_{i=1}^{n}$. Under interference, we are interested in the causal effect of $A^{s}$ on $Y$. Hence, the scientific question of interest is actually, “what if $A_{i}^{s}$ were replaced by some $A_{i}^{s\circ d}=s_{F_{i}}(d(\mathbf{A},\mathbf{L};\delta),\mathbf{L})$?” This process is illustrated in Figure 1. We call the function composition $s\circ d$ the induced MTP.
@@
Under an induced MTP, interference no longer hampers identifiability because $s_{F_{i}}(\mathbf{A},\mathbf{L})$ captures the contribution of all relevant units (that is, a given unit $i$ and its neighbors) to each $Y_{i}$. This data-adaptive parameter will converge to the population counterfactual mean as $n\rightarrow\infty$ (Hubbard et al., 2016). Use of such a parameter definition is necessary because we must condition on the single observation of the interference network at play. Hence, do note that in all theory that follows, our estimates will implicitly condition on the observed network $\mathbf{F}$. This estimand must be compatible with the network: it is interpreted as the average change in $Y$ caused by imposing the unit-level MTP $d$ on each unit in the population governed by the network $\mathbf{F}$. With $\Psi_{n}(\mathsf{P})$ identified, the population intervention effect (a contrast) may be defined by subtracting $\mathbb{E}Y$ (Díaz & van der Laan, 2012).
@@ 3.2 Identification
If the summaries satisfy $(s_{F_{i}}(\mathbf{a},\mathbf{l}),s_{F_{i}}(\mathbf{l}))\in\text{supp}\{A_{i}^{s},L_{i}^{s}\}$, then the summaries after intervention satisfy $(s_{F_{i}}(\mathbf{a}^{d},\mathbf{l}),s_{F_{i}}(\mathbf{l}))\in\text{supp}\{A_{i}^{s},L_{i}^{s}\}$.